# FOF-X: Towards Real-time Detailed Human Reconstruction from a Single Image

### v2

@@ III-A1 Vanilla FOF
An overview of our FOF representation is shown in Fig. 4. Without loss of generality, 3D objects to be reconstructed are normalized into a $[-1,1]^{3}$ cube. Denoted by $(x,y,z)\in[-1,1]^{3}$ the 3D coordinates, and by $S$ the surface of a 3D object, the 3D object can be described by a 3D occupancy field $F:[-1,1]^{3}\mapsto\{0,0.5,1\}$, where the occupancy function $F(x,y,z)$ is defined as
$F(x,y,z)=\left\{\begin{array}[]{ll}1,&\text{$(x,y,z)$ is inside the object},\\ 0.5,&\text{$(x,y,z)\in S$},\\ 0,&\text{$(x,y,z)$ is outside the object}.\end{array}\right.$ (1)
Such a 3D representation is highly redundant, because only a small subset, i.e., the iso-surface of $F(x,y,z)$ with value of 0.5, is sufficient to represent the surface of the object. Direct inference of the 3D occupancy field $F$ from a single image not only confronts with the curse of dimensionality, but also requires more computation and memory in handling high dimensional feature maps. Note that the occupancy field defined in the 3D cube can be regarded as the collection of a large number of 1D signals defined on the lines along the view direction. Without loss of generality, we assume that the view direction is the same as the $z$-axis of the 3D cube. The occupancy function, as a 1D signal $f(z):[-1,1]\mapsto\left\{0,0.5,1\right\}$, along the line passing through a particular point $(x^{*},y^{*})\in[-1,1]^{2}$ on the $xy$-plane can be written as $f(z)=F(x^{*},y^{*},z)$. We explore Fourier representation for such a family of 1D occupancy signals and propose a more compact 2D vector field orthogonal to the view direction, namely Fourier Occupancy Field (FOF), for more efficient representation of the 3D occupancy field, which is detailed in the following subsections.
@@
[…]
@@
[…]
Fourier Occupancy Field: Such a Fourier subspace approximation is applied to all 1D occupancy signals over the $xy$-plane:
@@
[…]
@@ III-A2 FOF with Cosine Series
Because $f(z)$ is defined on [1,1][-1,1][ - 1 , 1 ] but not [0,1][0,1][ 0 , 1 ], a shift on $z$ is needed, which contributes to the term $\frac{(z+1)}{2}$ in Eq. (5).
Figure 4: The proposed Fourier Occupancy Field (FOF). Each pixel on the image $I$ corresponds to a line in 3D space and can be described by an occupancy function $f(z)$, which can be expanded as Fourier series and represented with a vector $C(x,y)$. Therefore, the 3D human model can be encoded as a 2D vector field, forming our FOF. Note that the Fourier series here includes special forms such as the cosine series.
@@ III-B3 Supervision
To train our network, we sample points on the $xy$-plane and supervise the coefficients of the lines corresponding to these points with mean squared error (MSE) loss. To make the network more focused on human geometry, we only supervise the human foreground region of the image. The loss function $\mathcal{L}$ is formulated as:
@@ III-C1 Mesh to FOF with Automaton-based Discontinuity Matcher
where $t_{n}=\frac{n\pi}{2}$. The FOF representation is obtained by calculating the cosine coefficients for all the 1D occupancy signals. Since the number of intervals $k$ is generally small, e.g., 1$\sim$3 for human meshes, the calculation of $a_{n}$ is of $O(1)$ complexity. Thus, the transformation from discontinuities to FOF, shown in Alg. 1, has a computational complexity of $O(HWN)$, and can be parallelized on GPUs for real-time running, where $H$ and $W$ are the height and width of the FOF, respectively.
@@
1:A triangle mesh $M=\{t_{i}\}_{i=1}^{n}$triangleright A set of triangles
@@
13:      $B[x,y]\leftarrow B[x,y]\cup{(z,0)}$triangleright Add to buffer
@@
15:      $B[x,y]\leftarrow B[x,y]\cup{(z,1)}$triangleright Add to buffer
@@
However, due to numerical errors and some non-watertight meshes, rasterization may produce unpaired $\mathcal{Z}$, i.e., a line may encounter successive discontinuities both going out of or into the mesh. And it can come to a going outside point first instead of a going inside point. To solve this, we propose an automaton-based discontinuity matcher to enhance the robustness. Unlike what we do in the conference version [19], we collect those discontinuities and normal vectors for each line first and process them with an automaton before working out the integral. Let’s consider the discontinuities corresponding to one pixel. We sort these points by $z$ first and then represent them with a sequence of 0s and 111s, where 0 stands for the normal vector facing front and 111 stands for the normal vector facing backwards. We add a stop letter \$\\$\$ to the sequence and then process it with the automaton shown in Fig. 5, which effectively handles irregular patterns (initial 111s and repeating 0s or 111s). After being filtered by the automaton, the produced sequence can then be integrated with Eq. (8). For watertight meshes, both our automaton and the conference version produce identical results, correctly converting meshes to FOF. For non-watertight meshes (e.g., surfaces with open boundaries or self-intersections, partly missing, or with non-manifold connectivity), the conference version method fails to compute FOF accurately, leading to floating artifacts (Fig. 2). Our automaton-based discontinuity matcher robustly handles irregular patterns arising from non-watertight meshes (including holes, self-intersections, and duplicate faces), effectively preventing reconstruction artifacts caused by such topological imperfections. Fig. 6 shows an example of our automaton processing a non-watertight mesh.
@@ III-C2 FOF to Mesh with Laplacian Coordinate Constraint
where $D$ is a diagonal matrix with $D_{ii}=d_{i}$ on the diagonal, $d_{i}$ is the number of neighbors of the $i^{\textrm{th}}$ point, $X\in\mathbb{R}^{n\times 3}$ is the coordinate matrix to be optimized, and $A$ is the adjacency matrix. $A_{ij}=1$ if points $v_{i}$ and $v_{j}$ are adjacent, otherwise $A_{ij}=0$. $U$ is the index set of unreliable points, which are to be optimized. Unlike excessive Laplacian smoothing iterations that degrade geometric features into spherical approximations, our Laplacian coordinate constraint preserves topological fidelity and guarantees convergence to ground-truth mesh shapes.
@@ IV-B Implementation Details
The reconstruction network of FOF-X takes dual-sided normal maps and the FOF of SMPL as input. The dimensions of the input are 512×512×(3+3+16)512$\times$ 512$\times$(3+3+16)512 × 512 × ( 3 + 3 + 16 ), and the dimensions of the produced feature map are 128×128×256128$\times$ 128$\times$ 256128 × 128 × 256. We adopt HRNet-W48 [49] as the image encoder of our system due to its multi-scale perception ability. Since the SMPL mesh only provides a rough shape prior, the high-frequency components corresponding to geometric details are redundant for the input. Empirically, we find that the first 16 terms of the SMPL mesh is sufficient for the shape prior. The decoder progressively upsamples the feature map to $\times 256$ 256 and $\times 512$ 512 using bilinear interpolations and residual convolutional blocks, while retaining a channel number of 256256256 throughout the process. Finally, a convolutional layer projects the features to a FOF with a resolution of 512×512×128512$\times$ 512$\times$ 128512 × 512 × 128. To cope with single view image as input, we use the off-the-shelf image-to-image translation network used in ICON [3] for generating dual-sided normal maps from the single image.
@@
For consistency with previous works [1, 4, 3], we employ a weak perspective camera. The scale of the weak perspective camera is fixed at 111 for training and evaluation, resulting in an orthogonal projection. But note that we can transform world space to NDC (Normalized Device Coordinates) space to support perspective cameras straightforwardly. We follow the post-processing for hands described in ECON[4], fusing the hands of the parametric model with the reconstructed body. This strategy is not used in our real-time systems.
@@ IV-D Comparison with the State of the Arts
The quantitative results are summarized in Table II. denotes the re-implemented version in ICON[3]. denotes the re-implemented version with the cosine series and the improved inter-conversion algorithms. FOF-X(HR32) and FOF-X(HR48) are FOF-X model with HRNet-W32 and HRNet-W48 respectively. All methods listed, except PIFuHD[11] and PSHuman[43], are trained on THuman2.0[47]. It is worth noting that PSHuman’s[43] dataset includes THuman 2.1. For PIFu[1] and PaMIR[12], similar to ICON[3], we use the re-implemented versions, as the official releases are trained with additional commercial datasets. Since some methods do not provide the training sample lists, we use THuman2.1 instead of THuman2.0 to evaluate all methods. We use $\times 512\times 512$ 512 resolution for the input of all these methods. We do not compare our method with [13] and [51] because their codes are not released. For simplicity, we also ignore some works that have already been compared with the state-of-the-art methods listed above. We found that some methods are sensitive to the ratio of the human subject’s size in the image. To ensure consistency, we use a ratio of 0.9 for all methods, corresponding to the 1.8m height normalization mentioned in the datasets section. Ground-truth SMPL meshes are used to better demonstrate the capability of the evaluated methods.
@@ IV-E1 Number of Terms to Use
In the original FOF, we use $2N+1=31$ terms of the Fourier series to represent the occupancy field, which has been shown to be sufficient for maintaining geometric accuracy. However, we observe that the absence of high-frequency components can lead to visual artifacts. Therefore, we use $N=128$ terms of the cosine series in FOF-X. We illustrate the impact of varying $N$ on the accuracy of the approximation in Fig. 10. The metrics for different values of $N$ are also presented in Table III. When $N\leq 16$, noticeable artifacts appear on the meshes. Noting the geometric details in the ear region, although using $N=32$ preserves most geometric information, we prefer $N=128$ in FOF-X to ensure higher visual quality and eliminate potential artifacts.
@@ IV-E2 Sampling Scalability
The frequency-domain resolution of FOF can be adaptively inferred according to the time needs of the system, without the need for retraining. We explored the effect of resolution (number of components) on the results taking $N=128$. Fig. 11 illustrates the effect of changing resolution on the accuracy of the results. Metrics of different $N$ values are also given in Table IV.
@@ IV-E3 Noise Robustness
Figure 10: Results of different number $N$ of coefficients used in FOF-X.
@@
Figure 11: Results of different resolution $Res$ used in FOF-X. We use $Res=256$ in our implementation.
@@ IV-F The Real-time Pipeline
To perform real-time monocular human reconstruction, we design a four-stage pipeline. In the first stage, we get an image from the video stream with OpenCV [52], use RVM [53] to get a human mask to remove the background, and predict a SMPL model as prior with 4D-humans[54]. RVM is originally a matting method, and we use a threshold of 0.5 to obtain a binary mask. In the second stage, a CUDA-implemented rasterizer is used to render dual-sided normal maps and a FOF representation for the predicted SMPL. The proposed mesh-to-FOF conversion algorithm is integrated into the rasterizer. In the third stage, the pre-trained dual-sided normal estimator and our reconstruction network are encapsulated into a single neural network module, which is used to reconstruct the FOF. In the final stage, the FOF is resized to a proper resolution ($\times 256$ 256 in our implementation) and transformed into a mesh. All four stages are implemented in PyTorch and accelerated with TensorRT, running on a single RTX-4090 GPU. Our pipeline works at 30+ FPS, which is limited by the camera (Logic C920 Pro). With a recorded video, it can achieve a frame rate of about 50 FPS. Fig. 13 shows the real-time reconstruction system (all human subjects provided informed consent for data collection). Please check the demo video for more results.

### v3

@@ III-A1 Vanilla FOF
An overview of our FOF representation is shown in Fig. 4. Without loss of generality, 3D objects to be reconstructed are normalized into a $1,1]^{3}$ cube. Denoted by $(x,y,z)\in[-1,1]^{3}$ the 3D coordinates, and by $S$ the surface of a 3D object, the 3D object can be described by a 3D occupancy field $F:[-1,1]^{3}\mapsto\{0,0.5,1\}$, where the occupancy function $F(x,y,z)$ is defined as
[…]
Such a 3D representation is highly redundant, because only a small subset, i.e., the iso-surface of $F(x,y,z)$ with value of 0.5, is sufficient to represent the surface of the object. Direct inference of the 3D occupancy field $F$ from a single image not only confronts with the curse of dimensionality, but also requires more computation and memory in handling high dimensional feature maps. Note that the occupancy field defined in the 3D cube can be regarded as the collection of a large number of 1D signals defined on the lines along the view direction. Without loss of generality, we assume that the view direction is the same as the $z$-axis of the 3D cube. The occupancy function, as a 1D signal $f(z):[-1,1]\mapsto\left\{0,0.5,1\right\}$, along the line passing through a particular point $(x^{*},y^{*})\in[-1,1]^{2}$ on the xyxy-plane can be written as $f(z)=F(x^{*},y^{*},z)$. We explore Fourier representation for such a family of 1D occupancy signals and propose a more compact 2D vector field orthogonal to the view direction, namely Fourier Occupancy Field (FOF), for more efficient representation of the 3D occupancy field, which is detailed in the following subsections.
@@
where $\{a_{n}\},\{b_{n}\}\in\mathbb{R}$ are coefficients of basis functions $\left\{\cos(nx)\right\}$ and $\left\{\sin(nx)\right\}$, respectively. Note that by defining the occupancy function as Eq. (1), Eq. (2) holds for discontinuity at $z^{*}$ on the surface $S$ because $f(z^{*})=(f(z^{*}-)+f(z^{*}+))/2=0.5$, where $f(z^{*}-)$ and $f(z^{*}+)$ are the left hand limit and right hand limit of $f(z)$ at $z^{*}$. For compact representation, we approximate 1D occupancy signals by a subspace spanned by the first $2N+1$ basis functions:
@@
where $\bm{b}(z)=[1/2,\cos(z),\sin(z),\ldots,\cos(Nz),\sin(Nz)]^{\top}$ is the vector of the first $2N+1$ basis functions spanning the approximation subspace, and $\bm{c}=[a_{0},a_{1},b_{1},\ldots,a_{N},b_{N}]^{\top}$ is the coefficient vector which provides a more compact representation of the 1D occupancy function $f(z)$.
Fourier Occupancy Field: Such a Fourier subspace approximation is applied to all 1D occupancy signals over the xyxy-plane:
@@
where $\bm{C}(x,y)$ is the ($2N+1$)-dimensional Fourier coefficient vector for the 1D occupancy signal at $(x,y)$. In this way, we obtain the Fourier Occupancy Field $\bm{C}:[-1,1]^{2}\mapsto\mathbb{R}^{2N+1}$ for 3D occupancy field $F$.
@@ III-A2 FOF with Cosine Series
Because $f(z)$ is defined on [1,1][-1,1] but not $[0,1]$, a shift on $z$ is needed, which contributes to the term $\frac{(z+1)}{2}$ in Eq. (5).
Figure 4: The proposed Fourier Occupancy Field (FOF). Each pixel on the image II corresponds to a line in 3D space and can be described by an occupancy function $f(z)$, which can be expanded as Fourier series and represented with a vector $C(x,y)$. Therefore, the 3D human model can be encoded as a 2D vector field, forming our FOF. Note that the Fourier series here includes special forms such as the cosine series.
@@ III-B3 Supervision
To train our network, we sample points on the xyxy-plane and supervise the coefficients of the lines corresponding to these points with mean squared error (MSE) loss. To make the network more focused on human geometry, we only supervise the human foreground region of the image. The loss function $\mathcal{L}$ is formulated as:
@@ III-C1 Mesh to FOF with Automaton-based Discontinuity Matcher
1:A triangle mesh $M=\{t_{i}\}_{i=1}^{n}$ $\triangleright$ A set of triangles
@@
13:      $B[x,y]\leftarrow B[x,y]\cup{(z,0)}$ $\triangleright$ Add to buffer
@@
15:      $B[x,y]\leftarrow B[x,y]\cup{(z,1)}$ $\triangleright$ Add to buffer
@@
However, due to numerical errors and some non-watertight meshes, rasterization may produce unpaired $\mathcal{Z}$, i.e., a line may encounter successive discontinuities both going out of or into the mesh. And it can come to a going outside point first instead of a going inside point. To solve this, we propose an automaton-based discontinuity matcher to enhance the robustness. Unlike what we do in the conference version [19], we collect those discontinuities and normal vectors for each line first and process them with an automaton before working out the integral. Let’s consider the discontinuities corresponding to one pixel. We sort these points by $z$ first and then represent them with a sequence of 0s and 11s, where 0 stands for the normal vector facing front and 11 stands for the normal vector facing backwards. We add a stop letter \$\\$ to the sequence and then process it with the automaton shown in Fig. 5, which effectively handles irregular patterns (initial 11s and repeating 0s or 11s). After being filtered by the automaton, the produced sequence can then be integrated with Eq. (8). For watertight meshes, both our automaton and the conference version produce identical results, correctly converting meshes to FOF. For non-watertight meshes (e.g., surfaces with open boundaries or self-intersections, partly missing, or with non-manifold connectivity), the conference version method fails to compute FOF accurately, leading to floating artifacts (Fig. 2). Our automaton-based discontinuity matcher robustly handles irregular patterns arising from non-watertight meshes (including holes, self-intersections, and duplicate faces), effectively preventing reconstruction artifacts caused by such topological imperfections. Fig. 6 shows an example of our automaton processing a non-watertight mesh.
@@ III-C2 FOF to Mesh with Laplacian Coordinate Constraint
where $D$ is a diagonal matrix with $D_{ii}=d_{i}$ on the diagonal, $d_{i}$ is the number of neighbors of the $i^{\textrm{th}}$ point, $X\in\mathbb{R}^{n\times 3}$ is the coordinate matrix to be optimized, and AA is the adjacency matrix. $A_{ij}=1$ if points $v_{i}$ and $v_{j}$ are adjacent, otherwise $A_{ij}=0$. $U$ is the index set of unreliable points, which are to be optimized. Unlike excessive Laplacian smoothing iterations that degrade geometric features into spherical approximations, our Laplacian coordinate constraint preserves topological fidelity and guarantees convergence to ground-truth mesh shapes.
@@ IV-B Implementation Details
The reconstruction network of FOF-X takes dual-sided normal maps and the FOF of SMPL as input. The dimensions of the input are $512\times 512\times(3+3+16)$, and the dimensions of the produced feature map are $128\times 128\times 256$. We adopt HRNet-W48 [49] as the image encoder of our system due to its multi-scale perception ability. Since the SMPL mesh only provides a rough shape prior, the high-frequency components corresponding to geometric details are redundant for the input. Empirically, we find that the first 16 terms of the SMPL mesh is sufficient for the shape prior. The decoder progressively upsamples the feature map to $256\times 256$ and $512\times 512$ using bilinear interpolations and residual convolutional blocks, while retaining a channel number of 256256 throughout the process. Finally, a convolutional layer projects the features to a FOF with a resolution of $512\times 512\times 128$. To cope with single view image as input, we use the off-the-shelf image-to-image translation network used in ICON [3] for generating dual-sided normal maps from the single image.
@@
For consistency with previous works [1, 4, 3], we employ a weak perspective camera. The scale of the weak perspective camera is fixed at 11 for training and evaluation, resulting in an orthogonal projection. But note that we can transform world space to NDC (Normalized Device Coordinates) space to support perspective cameras straightforwardly. We follow the post-processing for hands described in ECON[4], fusing the hands of the parametric model with the reconstructed body. This strategy is not used in our real-time systems.
@@ IV-D Comparison with the State of the Arts
The quantitative results are summarized in Table II. denotes the re-implemented version in ICON[3]. denotes the re-implemented version with the cosine series and the improved inter-conversion algorithms. FOF-X(HR32) and FOF-X(HR48) are FOF-X model with HRNet-W32 and HRNet-W48 respectively. All methods listed, except PIFuHD[11] and PSHuman[43], are trained on THuman2.0[47]. It is worth noting that PSHuman’s[43] dataset includes THuman 2.1. For PIFu[1] and PaMIR[12], similar to ICON[3], we use the re-implemented versions, as the official releases are trained with additional commercial datasets. Since some methods do not provide the training sample lists, we use THuman2.1 instead of THuman2.0 to evaluate all methods. We use $512\times 512\times 512$ resolution for the input of all these methods. We do not compare our method with [13] and [51] because their codes are not released. For simplicity, we also ignore some works that have already been compared with the state-of-the-art methods listed above. We found that some methods are sensitive to the ratio of the human subject’s size in the image. To ensure consistency, we use a ratio of 0.9 for all methods, corresponding to the 1.8m height normalization mentioned in the datasets section. Ground-truth SMPL meshes are used to better demonstrate the capability of the evaluated methods.
@@ IV-E1 Number of Terms to Use
In the original FOF, we use $2N+1=31$ terms of the Fourier series to represent the occupancy field, which has been shown to be sufficient for maintaining geometric accuracy. However, we observe that the absence of high-frequency components can lead to visual artifacts. Therefore, we use $N=128$ terms of the cosine series in FOF-X. We illustrate the impact of varying NN on the accuracy of the approximation in Fig. 10. The metrics for different values of NN are also presented in Table III. When $N\leq 16$, noticeable artifacts appear on the meshes. Noting the geometric details in the ear region, although using $N=32$ preserves most geometric information, we prefer $N=128$ in FOF-X to ensure higher visual quality and eliminate potential artifacts.
@@ IV-E2 Sampling Scalability
The frequency-domain resolution of FOF can be adaptively inferred according to the time needs of the system, without the need for retraining. We explored the effect of resolution (number of components) on the results taking $N=128$. Fig. 11 illustrates the effect of changing resolution on the accuracy of the results. Metrics of different NN values are also given in Table IV.
@@ IV-E3 Noise Robustness
Figure 10: Results of different number NN of coefficients used in FOF-X.
@@
Figure 11: Results of different resolution ResRes used in FOF-X. We use $Res=256$ in our implementation.
@@ IV-F The Real-time Pipeline
To perform real-time monocular human reconstruction, we design a four-stage pipeline. In the first stage, we get an image from the video stream with OpenCV [52], use RVM [53] to get a human mask to remove the background, and predict a SMPL model as prior with 4D-humans[54]. RVM is originally a matting method, and we use a threshold of 0.5 to obtain a binary mask. In the second stage, a CUDA-implemented rasterizer is used to render dual-sided normal maps and a FOF representation for the predicted SMPL. The proposed mesh-to-FOF conversion algorithm is integrated into the rasterizer. In the third stage, the pre-trained dual-sided normal estimator and our reconstruction network are encapsulated into a single neural network module, which is used to reconstruct the FOF. In the final stage, the FOF is resized to a proper resolution ($256\times 256$ in our implementation) and transformed into a mesh. All four stages are implemented in PyTorch and accelerated with TensorRT, running on a single RTX-4090 GPU. Our pipeline works at 30+ FPS, which is limited by the camera (Logic C920 Pro). With a recorded video, it can achieve a frame rate of about 50 FPS. Fig. 13 shows the real-time reconstruction system (all human subjects provided informed consent for data collection). Please check the demo video for more results.